The units digit of a certain two digit number is one more than two times the tens digit. The number formed by interchanging the digits is 4 less than 2 times the original number. Find the original number. Please help me solve this!!!
Answer by mathstutor458(57)

let the number is xy
The units digit of a certain two digit number is one more than two times the tens digit.
so,y=2x+1
so,13,25,37,49 are the possible numbers.
first take xy=13
The number formed by interchanging the digits is 4 less than 2 times the original number.
which means, yx=2xy-4
=>31=2*13-4 does not satisfied
take xy=25
yx=2xy-4=>52=2*25-4 does not satisfied
take xy=37
yx=2xy-4=>73=2*37-4 does not satisfied
take xy=49
yx=2xy-4=>94=2*49-4=>94=94 satisfies the condition
Therefore the original number is 49